Air contained

Air contained in a rigid, insulated tank fitted with a paddle wheel, initially at 1 bar, 330 K and a volume of 1.93 m^3, receives an energy transfer by work from thepaddle wheel in an amount of 400kJ. Assuming the ideal gas model for the air, determine (a) the final temperature, in K, (b) the final pressure, in bar, and (c) theamount of entropy produced, in kJ/K, Ignore kinetic and potential energy.
Answer #1

As it is a rigid flask its a constant volume process

Cv of air is.717 KJ/kg K

pressure 1 bar is 10^5 pascals

work=Δinternal energy
400=n*Cv*ΔT
400=PV/RT *CvΔT
400 Kj= (1*10^5)*(1.93)/(8.314*(330)) *(.717 kj) (Tf - 330)
a)Tf=337.93 K
b)Pf*V=nRT
Pf/Tf=P/T
Pf/335.68 = 1/330
Pf=1.024 bar
c)Δentropy = n*Cv*loge(Tf/Ti) + R*loge(Vf/Vi) asVf=Vi --> ln(Vf/Vi) = 0
=50.4374*ln(335.68/330)=1.1977 kJ

answered by: Elleni